### v1

# Faithful global convergence for the rescaled Consensus–Based Optimization

Abstract.
We analyze the Consensus-Based Optimization (CBO) algorithm with a consensus point rescaled by a small fixed parameter $\kappa\in(0,1)$. Under minimal assumptions on the objective function and the initial data, we establish its unconditional convergence to the global minimizer. Our results hold in the asymptotic regime where both the time–horizon $t\to\infty$ and the inverse–temperature $\alpha\to\infty$, providing a rigorous theoretical foundation for the algorithm’s global convergence. Furthermore, our findings extend to the case of multiple and non–discrete set of minimizers.
Key words and phrases:
Consensus-Based Optimization; Laplace’s principle; global optimization
MSC Mathematics Subject Classification:
65C35, 90C26, 90C56
1. Introduction
Consensus-Based Optimization (CBO) is a derivative–free algorithm designed to tackle large–scale, highly non–convex optimization problems. This algorithm and its analytical foundations were established in [pinnau2017consensus, carrillo2018analytical]. It has since then been extensively studied, both theoretically and practically, demonstrating its effectiveness in solving complex optimization challenges that arise in real–world applications.
Inspired by the collective behavior observed in social and biological systems, CBO exploits the interactions of multiple agents, or particles, to explore the search space in a coordinated and efficient manner. The method uses a stochastic consensus mechanism, where particles are dynamically guided toward regions of lower objective values, ultimately converging to an optimal solution. This decentralized and adaptive search strategy makes CBO particularly well-suited for problems where gradient information is unavailable or unreliable. Motivated by its broad applicability, researchers have extended and refined the original CBO framework to accommodate various settings and demonstrated strong empirical success in solving wide range of global optimization problems [fornasier2021JMLR, ha2022stochastic, borghi2022consensus, herty2024multiscale, carrillo2022consensus, carrillo2021consensus, fornasier2022anisotropic, fornasier2020consensus, fornasier2024truncated, byeon2025consensus, chen2022consensus, wei2025consensus, bonandin2024consensus, cipriani2022zero]. These advancements have further enhanced the algorithm’s robustness and efficiency, making it a versatile tool across a wide range of scientific and engineering disciplines.
One of the key contributions of this manuscript is to establish its convergence, thus providing further theoretical validation of the method.
Let us consider the following optimization problem:
Find $\text{Find }\;x_{*}\in\operatorname*{argmin}_{x\in{\mathbb{R}}^{d}}f(x),\,$
where $f$ can be a non-convex non-smooth objective function that one wishes to minimize. Then the CBO dynamic considers the following system of $N$ interacting particles, denoted as $\{X_{\cdot}^{i}\}_{i=1}^{N}$, which satisfies
(1.1) ${\rm d}X_{t}^{i}=-\lambda(X_{t}^{i}-\mathfrak{m}_{\alpha}(\rho_{t}^{N}))\,{\rm d}t+\sigma D(X_{t}^{i}-\mathfrak{m}_{\alpha}(\rho_{t}^{N}))\,{\rm d}B_{t}^{i},\quad i=1,\dots,N=:[N]\,,$
where $\sigma>0$ is a real constant, $\rho_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i}}$ is the empirical measure associated to the particle system, and $\{B_{.}^{i}\}_{i=1}^{N}$ are $N$ independent $d$-dimensional Brownian motions. Here, we employ the anisotropic diffusion in the sense that $D(X):=\mbox{diag}(|X_{1}|,\dots,|X_{d}|)$ for any $X\in{\mathbb{R}}^{d}$, which has been proven to handle high-dimensional problems more effectively [carrillo2021consensus, fornasier2022anisotropic]. The current global consensus point $\mathfrak{m}_{\alpha}(\rho_{t}^{N})$ is meant to approximate the global minimum, and is defined with a weighted average of the particles as
(1.2) $\mathfrak{m}_{\alpha}(\rho_{t}^{N}):=\frac{\int_{{\mathbb{R}}^{d}}x\,\omega_{\alpha}^{f}(x)\;\rho_{t}^{N}({\rm d}x)}{\int_{{\mathbb{R}}^{d}}\omega_{\alpha}^{f}(x)\;\rho_{t}^{N}({\rm d}x)}\,,\quad\text{ where }\quad\omega_{\alpha}^{f}(x):=\exp(-\alpha f(x)).$
This choice of weight function $\omega_{\alpha}^{f}(\cdot)$ is motivated by the well-known Laplace’s principle [miller2006applied, MR2571413], see also [huang2024consensus, Appendix A.2]. Moreover, the particle system is initialized with independent and identically distributed (i.i.d.) data $\{X_{0}^{i}\}_{i=1}^{N}$, where each $X_{0}^{i}$ is distributed according to a given measure $\rho_{0}\in\mathscr{P}({\mathbb{R}}^{d})$ for all $i\in[N]$.
The convergence proof of the CBO method is typically conducted in the context of large-particle limit [gerber2023mean, huang2022mean]. Specifically, rather than analyzing the $N$–particle system (1.1) directly, one considers the limit as $N$ approaches infinity and examines the corresponding McKean–Vlasov process ${\overline{X}}_{\cdot}$, which is governed by the following equation:
(1.3) ${\rm d}{\overline{X}}_{t}=$ $-\lambda\left({\overline{X}}_{t}-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})\right)\,{\rm d}t\,+\,\sigma D\left({\overline{X}}_{t}-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})\right)\,{\rm d}B_{t}\,,$
$\text{with }\;\rho_{t}^{\alpha}\,:=\,\text{Law}({\overline{X}}_{t}).$
1.1. The convergence problem in the literature
There are numerous results in the literature on the convergence analysis of CBO methods, which can be broadly categorized into two different approaches. For instance, in [fornasier2024consensus], the authors employ $\mathbb{E}[|\overline{X}_{t}-x_{*}|^{2}]$ as a Lyapunov functional for (1.3). Their convergence result states that for any given accuracy $\varepsilon>0$, there exists a time $T_{*}>0$ depending on $\varepsilon$, such that for $\alpha>\alpha_{0}$ (also depending on $\varepsilon$), the following holds:
$\mathbb{E}\left[|\overline{X}_{t}-x_{*}|^{2}\right]\leq\mathbb{E}\left[|\overline{X}_{0}-x_{*}|^{2}\right]e^{-\theta t}$
for some constant $\theta>0$ and for all $t\in[0,T_{*}]$. Moreover, it holds that
$\mathbb{E}\left[|\overline{X}_{T_{*}}-x_{*}|^{2}\right]\leq\varepsilon.$
This approach has been further extended to establish the convergence of various CBO variants, including those incorporating general constraints [borghi2023constrained, beddrich2024constrained], multi-player games [chenchene2023consensus], min-max problems [borghi2024particle], multiple-minimizer problems [fornasier2025pde], clustered federated learning [carrillo2023fedcbo], and bi-level optimization [trillos2024cb].
However, a limitation of this method is that it does not allow for taking the limits $t\to\infty$ or $\alpha\to\infty$, meaning that convergence is guaranteed only within a neighborhood of $x_{*}$ with accuracy $\varepsilon$.
One possible approach to achieving convergence as $t\to\infty$ is through a variance-based method described in [carrillo2018analytical]. This approach involves defining the variance function
$V(t)=\mathbb{E}\left[|\overline{X}_{t}-\mathbb{E}[\overline{X}_{t}]|^{2}\right].$
The proof is then divided into two steps. First, it is shown that under well-prepared initial data and appropriately chosen parameters, for any fixed $\alpha$, the function $V(t)$ converges to zero exponentially fast as $t\to\infty$. This implies that $\overline{X}_{t}$ converges to some consensus point $x^{\alpha}$ as $t\to\infty$. Second, it is verified that the consensus point $x^{\alpha}$ can be made arbitrarily close (up to any given accuracy $\varepsilon$) to the global minimizer $x_{*}$ by choosing some fixed $\alpha$ that is sufficiently large. This argument has been further extended to establish the convergence of such as particle swarm optimization [huang2023global], CBO constrained on the sphere [fornasier2021JMLR], CBO for saddle point problems [huang2024consensus], and CBO with mirror descent [bungert2025mirrorcbo]. A similar idea is also employed in the convergence proof at the particle level [ko2022convergence].
However, similar to the first method, this approach does not allow taking the limit $\alpha\to\infty$ and it only guarantees convergence within an $\varepsilon$–neighborhood of the global minimizer.
In this paper, we will establish the first true convergence result for CBO method, demonstrating that both limits, $t\to\infty$ and $\alpha\to\infty$, can be taken. Up to our knowledge, this is the first result of this kind. The main idea relies on the following.
1.2. A key observation and the rescaled CBO
As it has been observed in [huang2024self], any Dirac measure is invariant for the dynamics (1.3) which makes it difficult to study its asymptotic properties. This non–uniqueness of an invariant probability measure is mainly due to two reasons: in the drift, the term $-X$ is not strong enough with respect to the mean–field term $\mathfrak{m}_{\alpha}(\cdot)$, and the diffusion degenerates. This is the main reason behind our modification of (1.3) whose aim is to address exactly these two mentioned issues, and which has already been proposed in [huang2024uniform, huang2024self, herty2024multiscale]. Hence, we consider a rescaled CBO given by
(1.4) ${\rm d}{\overline{X}}_{t}=$ $-\lambda\left({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})\right)\,{\rm d}t\,+\,\sigma\left(\delta\,\mathds{I}_{d}+D\left({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})\right)\right)\,{\rm d}B_{t}$
$\text{with }\;\rho^{\alpha}_{t}\,:=\,\text{Law}({\overline{X}}_{t}).$
complemented with an initial condition ${\overline{X}}_{0}\sim\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$, where $\mathds{I}_{d}$ is the $d$-dimensional identity matrix, $10<\kappa<$ 1 is a small positive constant, and $\delta>0$ is any positive constant. The name is justified by the fact that the contribution of the mean–field term $\mathfrak{m}_{\alpha}(\cdot)$ is re-scaled with a parameter $\kappa$ when compared to the contribution of $-{\overline{X}}$.
Introducing the parameter $\kappa$ makes all Dirac measures not invariant anymore, and intuitively, creates a disconnection between the limits in the time variable $t\to\infty$ and the inverse–temperature $\alpha\to\infty$, hence allowing to easily handle the convergence of the dynamics towards the (to–be) global minimum.
On the other hand, the positive parameter $\delta\,$ guarantees ellipticity (non–degeneracy) of the PDE satisfied by the invariant probability measure of the dynamics, and grants it nice properties. More precisely, we need $\delta>0$ in order to guarantee existence of an invariant probability measure as in [huang2024self, Proposition 3.4]. Uniqueness then follows using the empirical approximation in [huang2024self, Corollary 3.8] which holds with no positivity requirement on $\delta$. Its value however does not influence the asymptotic analysis that we shall perform in the sequel, or the results that we aim for.
Clearly, the underlying system of $N$ interacting particles is of the form
(1.5) ${\rm d}X_{t}^{i}=-\lambda\left(X_{t}^{i}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}^{N})\right){\rm d}t+\sigma\left(\delta\,\mathds{I}_{d}+D(X_{t}^{i}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}^{N})\right){\rm d}B_{t}^{i},\quad i\in[N]\,$
and $\rho_{t}^{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i}}$ is the empirical measure associated to the particle system. In the following Figure 1, using standard Euler–Maruyama scheme, we apply our rescaled CBO particle system (1.5) as described in [huang2024uniform, Algorithm 1] to the Rastrigin function $R(x):=10+(x-1)^{2}-10\cos(2\pi(x-1))\,,$ which is a popular benchmark nonconvex function in optimization with multiple local minimizers and a unique global minimizer.
Figure 1. We apply the rescaled CBO particle system (1.5) to the Rastrigin function $R(x)$, which has a unique global minimizer $x_{*}=1$ (the red star). The initial particles (the blue dots) are sampled uniformly in $[2,5]$. The simulation parameters are $N=2*10^{6},\lambda=1,\sigma=0.2,\alpha=10^{15},\delta=1,{\rm d}t=0.01,T=100$ and $\kappa=0.01$. The final output is $\frac{1}{\kappa}\frac{1}{N}\sum_{i\in[N]}X_{T}^{i}\approx\frac{1}{\kappa}{\mathbb{E}}[{\overline{X}}_{T}]$ (the green circle).
The advantage of the rescaled CBO (1.4) manifests already in several contexts:
  • In [herty2024multiscale], it ensures the validity of an averaging principle which reduces a singularly perturbed (or, multi–scaled) dynamics of the form (1.4), into an averaged one. This can be seen as a model reduction technique for solving multi–level optimization problems using multiple populations of interacting particles.
  • In [huang2024self], it allows to prove existence and uniqueness of an invariant probability measure $\rho_{*}^{\alpha}$, and to establish the long–time behavior of its law. Moreover, its approximation with empirical measures has been proven with an explicit rate of convergence, leading to the so-called Self–interacting CBO.
  • In [huang2024uniform], it helped to guarantee a uniform–in–time propagation of chaos for the population of particles evolving along the dynamics (1.5). This property is crucial for ensuring a stable and reliable long–term convergence, which secure the practical effectiveness of CBO methods.
The present manuscript is in the continuation of the former work [huang2024self, huang2024uniform], where the proof of global convergence is absent. Our main contributions are discussed next.
1.3. Highlights
The three main features of the present manuscript are as follows.
  • Asymptotic behavior when $t\to\infty$ and $\alpha\to\infty$: We obtain the first true convergence result in the literature in the sense that both limits, $t\to\infty$ and $\alpha\to\infty$, can be taken. More specifically, we establish the asymptotic behavior of the unique invariant ($t\to\infty$) probability measure $\rho_{*}^{\alpha}$ of (1.4) as $\alpha\to\infty$. More precisely, when the function $f(\cdot)$ to be minimized has a unique global minimizer $x_{*}$, and for a given fixed parameter $\kappa$, we prove that
    $\lim_{\alpha\to\infty}\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\lim_{\alpha\to\infty}\int_{\mathbb{R}^{d}}x\,\rho_{*}^{\alpha}(\mbox{\rm d}x)=\kappa\,\lim_{\alpha\to\infty}\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})=\kappa\,x_{*}\,,$
    where ${\overline{X}}_{\cdot}$ satisfies (1.4).
  • Unconditional convergence: We are able to bypass classical assumptions in the literature.
    • Our results hold without any prior assumptions on the initial distribution $\rho_{0}$ of (1.4) or (1.5), provided it has sufficiently many finite moments. In particular, we do not assume that a global minimizer of $f(\cdot)$ is contained within the support of $\rho_{0}$, a crucial assumption in both [carrillo2018analytical] and [fornasier2024consensus]. Very recently, [fornasier2025regularity] demonstrated that this assumption can actually be removed by instead requiring regular initial data $\rho_{0}\in H^{m}(\mathbb{R}^{d})$. In contrast, our approach does not impose any regularity conditions on $\rho_{0}$, thereby allowing cases where $\rho_{0}$ is a Dirac measure, as has been considered in the self–interacting CBO model [huang2024self].
    • No regularity assumptions on the function $f(\cdot)$ are needed, besides being positive, locally Lipschitz continuous, and with a quadratic growth (see Assumption 2.1). These assumptions were used in [carrillo2018analytical] to guarantee the well-posedness of the CBO dynamics (1.3). To establish global convergence, [carrillo2018analytical, Assumption 4.1] imposes additional conditions, such as $f\in\mathcal{C}^{2}(\mathbb{R}^{d})$ and the boundedness of $\|\nabla^{2}f\|_{\infty}$. In the framework of [fornasier2024consensus, Definition 8], some additional local growth conditions on $f$ around the global minimizer are required. For our rescaled CBO we do not require these additional assumptions for the convergence proof.
  • Multiple–minimizer case: Although our rescaled CBO is not specifically designed to address the multiple minimizer problem, we can still establish certain asymptotic properties in cases where the objective function $f(\cdot)$ has multiple minimizers, whether they are discrete points or compact subsets (see Theorem 3.9). Let us mention here that specially designed CBO models aimed at addressing multiple-minimizer problems have been introduced in [fornasier2025pde, bungert2022polarized].
Our results rely on two main ingredients: uniform estimates on the moments of the measure, and on a Harnack inequality on the invariant elliptic CBO equation. Indeed we need Harnack’s inequality in order to ensure that the measure has a non–null mass around minimizers, which will then allow us to prove Laplace’s principle. A summary of our findings is in section 4 at the end of the manuscript.
2. Useful estimates
We start by recalling some notation. When a measure $\mu$ has a density $\varrho$ with respect to Lebesgue measure that we denote by $\mbox{\rm d}x$, then $\mu$ is absolutely continuous with respect to $\text{d}x$, we write $\mu\ll\mbox{\rm d}x$ and $\varrho=\frac{\text{d}\mu}{\text{d}x}$ is Radon–Nikodym derivative of $\mu$ with respect to $\mbox{\rm d}x$. Unless a confusion arises, we shall use the same notation for a measure and for its density (when it exists). We shall also denote by $W^{k,p}(\Omega)$, $p\geq 1,k\geq 0$ the standard Sobolev space of functions whose generalized derivatives up to order $k$ are in $L^{p}(\Omega)$, and we denote by $W^{k,p}_{loc}(\Omega)$ the class of functions $f$ such that $\chi f\in W^{k,p}(\Omega)$ for each $\chi\in\mathcal{C}_{c}^{\infty}(\Omega)$ the class of infinitely differentiable functions with compact support in $\Omega$.
[…]
Our main assumption concerns the function $f(\cdot)$ to be minimized, and can be stated as follows.
Assumption 2.1.
We assume the following properties for the objective function.
  1. (1)
    $f:~{}{\mathbb{R}}^{d}\to{\mathbb{R}}_{+}$ is bounded from below by $\underline{f}:=\min f>0$, and there exists some constant $L_{f}>0$ such that
    $|f(x)-f(y)|\leq L_{f}(1+|x|+|y|)|x-y|,\quad\forall x,y\in{\mathbb{R}}^{d}\,.$
  2. (2)
    There exist constants $c_{\ell},c_{u}>0$ and $M>0$ such that
    $f-\underline{f}\leq c_{u}(|x|^{2}+1),\quad\forall x\in{\mathbb{R}}^{d}\;\mbox{and }\;c_{\ell}|x|^{2}\leq f-\underline{f}\mbox{ for }|x|\geq M\,.$
As we have mentioned earlier in the introduction, our main ingredients are moments bounds and a Harnack inequality satisfied by the invariant probability measure. These are the object of the current section.
2.1. Moments bounds
The mean–field term $\mathfrak{m}_{\alpha}(\mu)$ plays a crucial role in the behavior of the dynamics. Its properties are essential for our analysis to be carried out. Thus we shall recall some estimates on $\mathfrak{m}_{\alpha}(\mu)$ that we borrow from [carrillo2018analytical, Lemma 3.3].
Lemma 2.2.
Suppose that $f(\cdot)$ satisfies Assumption 2.1, then it holds that
(2.1) $|\mathfrak{m}_{\alpha}(\nu)|^{2}\leq\frac{\int_{{\mathbb{R}}^{d}}|x|^{2}e^{-\alpha f(x)}\nu(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(x)}\nu(\mbox{\rm d}x)}\leq b_{1}+b_{2}\int_{{\mathbb{R}}^{d}}|x|^{2}\mbox{\rm d}\nu\quad\forall\nu\in\mathscr{P}_{2}(\mathbb{R}^{d})\,,$
where
$b_{1}=M^{2}+b_{2}\quad\text{ and }\quad b_{2}=2\,\frac{c_{u}}{c_{\ell}}\left(1+\frac{1}{\alpha\,M^{2}\,c_{\ell}}\right).$
Remark 2.3.
In fact, as it has been already observed in [carrillo2018analytical, Remark .1], the constant $b_{2}$ can be chosen to be independent of $\alpha$ for $\alpha\geq 1$.
Lemma 2.4.
Let $f$ satisfy Assumption 2.1 and $\rho_{0}\in\mathscr{P}_{p}({\mathbb{R}}^{d})$ for any $p\geq 2$, then both the dynamics (1.5) and (1.4) have unique strong solutions and they satisfy that
[…]
Proof.
The proof of well–posedness for the particle system (1.5) follows the same reasoning as in [carrillo2018analytical, Theorem 2.1] or [gerber2023mean, Theorem 2.2], relying on the fact that the coefficients in (1.5) are locally Lipschitz. For the mean-field dynamics (1.4), well-posedness can be established using the Leray-Schauder fixed-point theorem, as demonstrated in [carrillo2018analytical, Theorem 3.1] or [gerber2023mean, Theorem 2.2]. Finally, the moment estimates in (2.2) are derived from [carrillo2018analytical, Lemma 3.4] or [gerber2023mean, Lemma 3.5] by using the Burkholder–Davis–Gundy inequality. ∎
We can now prove the following uniform–in–time second moment bound for ${\overline{X}}_{t}$ which will play a key role in the sequel.
Theorem 2.5.
Let $f(\cdot)$ satisfy assumption 2.1 and it satisfies $\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$. Consider $({\overline{X}}_{t})_{t\geq 0}$ with $\rho_{0}$-distributed initial data satisfying (1.4). Then for sufficiently small $\kappa$ there exists some constant $C_{2}$ depending only $\lambda,\sigma,d,\delta,\kappa$ and $f$ such that it holds
(2.3) $\sup_{t\geq 0}{\mathbb{E}}[|{\overline{X}}_{t}|^{2}]\leq\mathbb{E}\left[|{\overline{X}}_{0}|^{2}\right]+C_{2}\,.$
Proof.
Let $h(x):=\frac{1}{2}|x|^{2}$, for $x\in\mathbb{R}^{d}$. Then applying Itô’s formula leads to
${\rm d}h({\overline{X}}_{t})$ =$=\bigg{\{}-\lambda\nabla h({\overline{X}}_{t})\cdot\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big{)}$
$\quad\quad\quad+\frac{\sigma^{2}}{2}\text{trace}\left[D^{2}h({\overline{X}}_{t})\bigg{(}\delta\,\mathds{I}_{d}+D\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big{)}\bigg{)}^{2}\right]\bigg{\}}\,{\rm d}t$
$\quad\quad\quad\quad\quad\quad+\sigma\,\nabla h({\overline{X}}_{t})^{\top}\bigg{(}\delta\,\mathds{I}_{d}+D\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big{)}\bigg{)}\,{\rm d}B_{t}$
$=:F(t,{\overline{X}}_{t})\,{\rm d}t+G(t,{\overline{X}}_{t})\,{\rm d}B_{t}.$
One could observe that, thanks to Lemma 2.4 we have
$\mathbb{E}\left[\int_{0}^{t}G(s,{\overline{X}}_{s})\,{\rm d}B_{s}\right]=0,\quad\forall\,t>0.$
This is true in particular because of the well-posedness of the process ${\overline{X}}_{\cdot}$ according to Lemma 2.4 with $X_{0}^{i}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$, and its finite moments which guarantee that $\mathbb{E}\left[\int_{0}^{t}|G(s,{\overline{X}}_{s})|^{2}{\rm d}s\right]\leq C\,{\mathbb{E}}\left[\int_{0}^{t}|{\overline{X}}_{s}|^{4}{\rm d}s\right]<+\infty$. Therefore we have
(2.4) $\frac{{\rm d}}{{\rm d}t}\mathbb{E}\left[\frac{1}{2}|{\overline{X}}_{t}|^{2}\right]=\mathbb{E}[F(t,{\overline{X}}_{t})],\quad\forall\,t>0.$
We can now further estimate the right-hand side as follows. Recall
$F(t,{\overline{X}}_{t})$ $=-\lambda\nabla h({\overline{X}}_{t})\cdot\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{N}_{t})\big{)}$
$\quad+\frac{\sigma^{2}}{2}\text{trace}\left[D^{2}h({\overline{X}}_{t})\bigg{(}\delta\,\mathds{I}_{d}+D\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{N}_{t})\big{)}\bigg{)}^{2}\right]$
and with $h(x)=\frac{1}{2}|x|^{2}$ we have $\nabla h(x)=x$, and $D^{2}h(x)=\mathds{I}_{d}$. We deal first with the trace term in $F(t,{\overline{X}}_{t})$. We have
$\text{trace}\left[D^{2}h({\overline{X}}_{t})\bigg{(}\delta\,\mathds{I}_{d}+D\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big{)}\bigg{)}^{2}\right]$
$\quad\quad=\;\text{trace}\left[D^{2}h({\overline{X}}_{t})\bigg{(}\delta^{2}\,\mathds{I}_{d}+D\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big{)}^{2}+2\delta\,D\big{(}{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})\big{)}\bigg{)}\right]$
$\quad\quad=\;\delta^{2}\,\text{trace}(D^{2}h({\overline{X}}_{t}))+\text{trace}\big{[}D^{2}h({\overline{X}}_{t})D({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}))^{2}\big{]}$
$\quad\quad\quad\quad+2\delta\,\text{trace}\big{[}D^{2}h({\overline{X}}_{t})D({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}))\big{]}\;$
$\quad\quad\leq\;\delta^{2}d+\sum\limits_{k=1}^{d}|({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}))_{k}|^{2}+2\delta\sum\limits_{k=1}^{d}|({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}))_{k}|$
$\quad\quad\leq\;\delta^{2}d+\delta d+(\delta+1)|{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})|^{2}.$
In the last line, we used $|({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}))_{k}|\leq\frac{1}{2}(|({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}))_{k}|^{2}+1)$, then the sum yields $\sum\limits_{k=1}^{d}|({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}))_{k}|\leq\frac{1}{2}(|{\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t})|^{2}+d)$. Therefore we have
$F(t,{\overline{X}}_{t})$ $\leq-\lambda\left(1-\frac{\kappa}{2}\right)|{\overline{X}}_{t}|^{2}+\lambda\,\frac{\kappa}{2}|\mathfrak{m}_{\alpha}(\rho_{t})|^{2}$
$\quad\quad\frac{\sigma^{2}}{2}\delta(\delta+1)d+\sigma^{2}(\delta+1)|{\overline{X}}_{t}|^{2}+\sigma^{2}(\delta+1)\kappa^{2}|\mathfrak{m}_{\alpha}(\rho_{t})|^{2}$
$\leq\left\{-\lambda\left(1-\frac{\kappa}{2}\right)+\sigma^{2}\left(1+\delta\right)\right\}\,|{\overline{X}}_{t}|^{2}$
+$\quad\quad+\left\{\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right\}\,|\mathfrak{m}_{\alpha}(\rho_{t})|^{2}+\frac{\sigma^{2}}{2}\delta(\delta+1)d.$
Recalling (2.1), it holds that
$|\mathfrak{m}_{\alpha}(\rho_{t})|^{2}\leq b_{1}+b_{2}\,{\mathbb{E}}[|{\overline{X}}_{t}|^{2}]\,,$
which implies that
${\mathbb{E}}[F(t,{\overline{X}}_{t})]$$\leq\left\{-\lambda\left(1-\frac{\kappa}{2}\right)+\sigma^{2}\left(1+\delta\right)+b_{2}\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)\right\}\,{\mathbb{E}}[|{\overline{X}}_{t}|^{2}]$
$\quad\quad+\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)b_{1}+\frac{\sigma^{2}}{2}\delta(\delta+1)d.$
The latter, together with (2.4) finally yields
$\frac{{\rm d}}{{\rm d}t}\mathbb{E}\left[\frac{1}{2}|{\overline{X}}_{t}|^{2}\right]$ $=\mathbb{E}\left[F(t,{\overline{X}}_{t})\right]\leq-\gamma\,\mathbb{E}\left[|{\overline{X}}_{t}|^{2}\right]+C_{2}$
where
(2.5) $\gamma:=\lambda\left(1-\frac{\kappa}{2}\right)-\sigma^{2}\left(1+\delta\right)-b_{2}\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)$
$\text{and }\quad C_{2}:=\left(\lambda\,\frac{\kappa}{2}+\sigma^{2}\left(1+\delta\right)\kappa^{2}\right)b_{1}+\frac{\sigma^{2}}{2}\delta(\delta+1)d.$
It suffices to choose $\lambda,\sigma,\delta$, and $\kappa$ in order to guarantee $\gamma>0$, which together with Grönwall’s inequality yields
(2.6) $\mathbb{E}\left[|{\overline{X}}_{t}|^{2}\right]\leq\mathbb{E}\left[|{\overline{X}}_{0}|^{2}\right]\,e^{-\gamma\,t}+C_{2}\,\leq\mathbb{E}\left[|{\overline{X}}_{0}|^{2}\right]+C_{2}.$
Remark 2.6.
A configuration of the parameters where (2.3) is guaranteed can be obtained as follows. Recalling the notation in (2.5)-(2.6), we can choose $\delta=1$, then we need to satisfy the inequality
$\gamma=\lambda\left(1-\frac{\kappa}{2}\right)-2\,\sigma^{2}-b_{2}\left(\lambda\,\frac{\kappa}{2}+2\sigma^{2}\kappa^{2}\right)\;>0$
[…]
Choosing $\kappa<\frac{1}{2}(1+b_{2})^{-1}$ yields in particular $(1-\,\kappa\,(1+b_{2}))>\frac{1}{2}$ and $b_{2}\,\kappa^{2}<1$. Therefore, we have
$\lambda\;(1-\,\kappa\,(1+b_{2}))>\frac{\lambda}{2}\quad\text{ and}\quad 8\,\sigma^{2}>4\,\sigma^{2}(1+b_{2}\kappa^{2}).$
Conclusion: to guarantee $\gamma>0$, it suffices to choose for example
$\delta=1,\quad\kappa<\frac{1}{2}(1+b_{2})^{-1},\quad\text{ and }\quad\lambda>16\,\sigma^{2}.$
The next result captures the asymptotic behavior when $t\to\infty$ in (2.3), and concludes the first step of the desired global convergence of the rescaled CBO to the global minimum.
Lemma 2.7.
Let $\rho_{*}^{\alpha}\in\mathscr{P}_{2}({\mathbb{R}}^{d})$ be the unique invariant probability measure to the rescaled CBO (1.4). Then it holds that
(2.7) $\rho_{*}^{\alpha}(|\cdot|^{2})\leq\rho_{0}(|\cdot|^{2})+C_{2}\,,$
where $C_{2}$ comes from Theorem 2.5
Proof.
The existence has been proven in [huang2024self, Proposition 3.4]. Now we show that (2.7) holds. According to [zhang2023existence, Lemma 2.7], for any $|\varphi(x)|\leq C(1+|x|)$ it holds that
(2.8) $\lim_{t\to\infty}\int_{{\mathbb{R}}^{d}}\varphi(x)\,\mbox{\rm d}\rho_{t}^{\alpha}(x)=\int_{{\mathbb{R}}^{d}}\varphi(x)\,\mbox{\rm d}\rho_{*}^{\alpha}(x)\,.$
Indeed, for the latter result to hold, one needs [zhang2023existence, assumption (H1) with $r_{1}=1$] whose validity in our context has already been verified in [huang2024self, Appendix A.2].
Subsequently, for any fixed $r>0$, we can construct a $\mathcal{C}_{c}^{\infty}({\mathbb{R}}^{d})$ bump function $\phi_{r}$ such that it has support on the closed ball $B_{2r}(0)$ satisfying $|\phi_{r}(x)|\leq 1$ and is equal to 1 in $B_{r}(0)$. Then one can use (2.8) with $\varphi(x)=\phi_{r}(x)|x|^{2}$, and obtain
$\lim_{t\to\infty}\int_{{\mathbb{R}}^{d}}\phi_{r}(x)|x|^{2}\,\mbox{\rm d}\rho_{t}^{\alpha}(x)=\int_{{\mathbb{R}}^{d}}\phi_{r}(x)|x|^{2}\,\mbox{\rm d}\rho_{*}^{\alpha}(x)\,\mbox{ for any }r>0\,.$
Recalling the uniform-in-time estimate in (2.3), we have
$\int_{{\mathbb{R}}^{d}}\phi_{r}(x)|x|^{2}\,\mbox{\rm d}\rho_{*}^{\alpha}(x)=\lim_{t\to\infty}\int_{{\mathbb{R}}^{d}}\phi_{r}(x)|x|^{2}\,\mbox{\rm d}\rho_{t}^{\alpha}(x)\leq\limsup_{t\to\infty}\int_{{\mathbb{R}}^{d}}|x|^{2}\,\mbox{\rm d}\rho_{t}^{\alpha}(x)\leq\rho_{0}(|\cdot|^{2})+C_{2}\,.$
Furthermore, one can use Fatou’s lemma and obtain
$\rho_{*}^{\alpha}(|\cdot|^{2})=\int_{{\mathbb{R}}^{d}}|x|^{2}\,\mbox{\rm d}\rho_{*}^{\alpha}(x)\leq\liminf_{r\to\infty}\int_{{\mathbb{R}}^{d}}\phi_{r}(x)|x|^{2}\,\mbox{\rm d}\rho_{*}^{\alpha}(x)\leq\rho_{0}(|\cdot|^{2})+C_{2}\,.$
2.2. Harnack’s inequality
The result in (2.7) together with (2.1) ensure that $\mathfrak{m}_{\alpha}(\cdot)$ is well-defined (finite) when evaluated in the invariant probability measure $\rho_{*}^{\alpha}$. Namely, it holds that
(2.9) $|\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})|\leq(b_{1}+b_{2}\rho_{*}^{\alpha}(|\cdot|^{2}))^{\frac{1}{2}}\leq(b_{1}+b_{2}(\rho_{0}(|\cdot|^{2})+C_{2}))^{\frac{1}{2}}<\infty\,.$
Therefore, we can introduce the following matrix-valued and vector-valued functions defined for $(x,\mu)\in{\mathbb{R}}^{d}\times\mathscr{P}_{2}({\mathbb{R}}^{d})$ by
$A:=(a^{ij})_{i,j},\quad\text{ such that }\quad a^{ij}(x,\mu):=\frac{\sigma^{2}}{2}(\delta+|(x-\kappa\,\mathfrak{m}_{\alpha}(\mu))_{i}|)^{2}\delta_{ij}$
and $b:=(b^{i})_{i},\quad\text{ such that }\quad b^{i}(x,\mu):=-\lambda(x-\kappa\,\mathfrak{m}_{\alpha}(\mu))_{i}$
$\delta_{ij}$ being the Kronecker symbol (not to be confused with the positive constant $\delta$). In particular, when evaluated in the invariant probability measure $\rho_{*}^{\alpha}$, one obtains smooth coefficients (in $\mathcal{C}^{\infty}({\mathbb{R}}^{d})$)
$A_{*}:=(a^{ij}_{*})_{i,j},\quad\text{ such that }\quad a^{ij}_{*}(x):=a^{ij}(x,\rho_{*}^{\alpha})$
and $b_{*}:=(b^{i}_{*})_{i},\quad\text{ such that }\quad b^{i}_{*}(x):=b^{i}(x,\rho_{*}^{\alpha}).$
Moreover, the matrix $A_{*}$ is positive–definite. This identification is possible since the dependence of $A$ and $b$ on the measure is non-local – that is, it involves a quantity of the form $\int_{{\mathbb{R}}^{d}}K(y)\mbox{\rm d}\mu(y)$ which is moreover constant in $x$ (since the kernel $K$ does not involve $x$)–, rather than a pointwise (local) evaluation of (the density of) the measure $\mu$. In other words, regularity of (the density of) the measure $\mu$ does not influence regularity of the coefficients $A$ and $b$.
More generally, given a measure $\nu\in\mathscr{P}_{2}({\mathbb{R}}^{d})$, the quantity $|\mathfrak{m}_{\alpha}(\nu)|$ is well-defined (finite) thanks to (2.1), and we can define the following functions on ${\mathbb{R}}^{d}$
$A_{\nu}:=(a^{ij}_{\nu})_{i,j},\quad\text{ such that }\quad a^{ij}_{\nu}(x):=a^{ij}(x,\nu)$
and $b_{\nu}:=(b^{i}_{\nu})_{i},\quad\text{ such that }\quad b^{i}_{\nu}(x):=b^{i}(x,\nu).$
The diffusion (linear) operator whose coefficients are $A_{\nu}$ and $b_{\nu}$ is given by
$L_{A,b,\nu}\psi(x)=\text{trace}\big{(}A_{\nu}(x)D^{2}\psi(x)\big{)}+b_{\nu}(x)\cdot\nabla\psi(x)\quad\forall\;\psi\in\mathcal{C}_{c}^{\infty}({\mathbb{R}}^{d}),$
Its adjoint is denoted by $L_{A,b,\nu}^{*}$, and a measure $\mu$ is solution to $L_{A,b,\nu}^{*}\mu=0$ if it is a solution in the distributional sense that is
$\int_{{\mathbb{R}}^{d}}L_{A,b,\nu}\psi(x)\;\mbox{\rm d}\mu(x)=0\quad\forall\;\psi\in\mathcal{C}_{c}^{\infty}({\mathbb{R}}^{d}).$
The diffusion (linear) operator whose coefficients are $A_{*}$ and $b_{*}$ will be denoted by $L_{A,b,*}$, and its adjoint is $L_{A,b,*}^{*}$. It should be noted that the choice of the test functions being $\psi\in\mathcal{C}_{c}^{\infty}({\mathbb{R}}^{d})$ can be relaxed to include a larger space of functions but which we shall not do here for the sake of simplicity.
When we allow the coefficients $A$ and $b$ to be also functions of the measure argument, the resulting diffusion operator is non-linear. Analogously, we say that $\mu$ is solution to $L_{A,b,\mu}^{*}\mu=0$ if it is a solution in the distributional sense that is
$\int_{{\mathbb{R}}^{d}}L_{A,b,\mu}\psi(x)\;\mbox{\rm d}\mu(x)=0\quad\forall\;\psi\in\mathcal{C}_{c}^{\infty}({\mathbb{R}}^{d}).$
The following result is well-known and can be found for example in [bogachev2022fokker].
Theorem 2.8.
Let $\nu\in\mathscr{P}_{2}({\mathbb{R}}^{d})$ be fixed. Let $A_{\nu}$ and $b_{\nu}$ be as defined above. Let $\mu$ be a locally finite Borel measure satisfying $L_{A,b,\nu}^{*}\mu=0$. Then $\mu$ has a density denoted by $\varrho$ in $W_{\text{loc}}^{1,p}({\mathbb{R}}^{d})$ with $(p>d)$ that is locally Hölder continuous. If moreover $\mu$ is nonnegative, then the continuous version of the density $\varrho$ of $\mu$ satisfies Harnack’s inequality: for every compact set $K$ contained in a connected set $U$, it holds
$\sup_{K}\varrho\leq C\inf_{K}\varrho$
where $C$ depends only on $\|a^{ij}_{\nu}\|_{W^{1,p}(U)}$, $\|b^{i}_{\nu}\|_{L^{p}(U)}$, $\inf_{U}\det(A_{\nu})$, and $K$.
Proof.
First observe that the coefficients $A_{\nu}$ and $b_{\nu}$ are smooth, and $A_{\nu}$ is positive–definite. Then, existence and regularity of the density is [bogachev2022fokker, Corollary 1.6.7]. Harnack’s inequality is [bogachev2022fokker, Corollary 1.7.2]. ∎
Corollary 2.9.
The unique invariant probability measure $\rho_{*}^{\alpha}$ to the rescaled CBO (1.4) has a density (still denoted by $\rho_{*}^{\alpha}$) in $W_{\text{loc}}^{1,p}({\mathbb{R}}^{d})$ with $(p>d)$ that is locally Hölder continuous. Moreover, its continuous version satisfies Harnack’s inequality: for any ball $B_{r}(x_{0}):=\big{\{}x\in{\mathbb{R}}^{d}:|x-x_{0}|\leq r\big{\}}$ with $x_{0}\in{\mathbb{R}}^{d}$, one has
(2.10) $\sup_{x\in B_{r}(x_{0})}\rho_{*}^{\alpha}(x)\leq C\inf_{x\in B_{r}(x_{0})}\rho_{*}^{\alpha}(x)$
where $C$ depends only on $\|a^{ij}\|_{W^{1,p}(B_{2r}(x_{0}))}$, $\|b^{i}\|_{L^{p}(B_{2r}(x_{0}))}$ and $\inf_{B_{2r}(x_{0})}\det(A)$.
Proof.
Existence and uniqueness of an invariant probability measure $\rho_{*}^{\alpha}$ is obtained in [huang2024self, Proposition 3.4 & Corollary 3.8], that is, $\rho_{*}^{\alpha}$ is the unique solution to the (non-linear) equation $L_{A,b,\mu}^{*}\mu=0$ in the class of probability measures. Moreover, (2.7) ensures that $\rho_{*}^{\alpha}$ has a finite second moment. Therefore, we can consider the diffusion (linear) operator $L_{A,b,*}$ and we are in the setting of Theorem 2.8 which gives the desired result when we choose the solution to $L_{A,b,*}^{*}\mu=0$ within the probability measures. Indeed in this case, the unique probability measure $\mu$ satisfying $L_{A,b,*}^{*}\mu=0$ is $\rho_{*}^{\alpha}$. Finally, the sets $U$ and $K$ in Theorem 2.8 are chosen as the closed balls $B_{2r}(x_{0})$ and $B_{r}(x_{0})$ respectively. ∎
Lemma 2.10.
Assume $f(\cdot)$ satisfies Assumption 2.1, and let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.4). Let $x_{\circ}\in{\mathbb{R}}^{d}$ be an arbitrarily fixed point, then for any $\varepsilon>0$ there exists some $C_{\varepsilon}>0$ depending only on $\rho_{0}(|\cdot|^{2}),\,C_{2},\,x_{\circ}$ and $\varepsilon$ such that
(2.11) $\rho_{*}^{\alpha}(B_{\varepsilon}(x_{\circ}))\geq C_{\varepsilon},$
with the constant $C_{2}$ being the one from Theorem 2.5.
Proof.
For any $\varepsilon>0$, let us consider some $R>\varepsilon$ and use Harnack’s inequality (2.10)
$\sup\limits_{B_{R}(x_{\circ})}\rho_{*}^{\alpha}$ $\leq C_{B_{2R}(x_{\circ})}\inf\limits_{B_{R}(x_{\circ})}\rho_{*}^{\alpha}$
$\leq C_{B_{2R}(x_{\circ})}\inf\limits_{B_{\varepsilon}(x_{\circ})}\rho_{*}^{\alpha}\leq C_{B_{2R}(x_{\circ})}\frac{1}{|B_{\varepsilon}(x_{\circ})|}\rho_{*}^{\alpha}(B_{\varepsilon}(x_{\circ}))\,.$
Here $C_{B_{2R}(x_{\circ})}$ is a positive constant depending on $\|a^{ij}\|_{W^{1,p}(B_{2R}(x_{0}))}$, $\|b^{i}\|_{L^{p}(B_{2R}(x_{0}))}$ and $\inf_{B_{2R}(x_{0})}\det(A)$. Using (2.9), it is not difficult to see that these quantities depend only on $B_{2R}(x_{\circ})$ and on the constants $\delta,\kappa,\lambda,\sigma,R,b_{1},b_{2},C_{2}$ which are independent from $\alpha$.
On the other hand, it is easy to check that
$\sup\limits_{B_{R}(x_{\circ})}\rho_{*}^{\alpha}$ $\geq\frac{1}{|B_{R}(x_{\circ})|}\rho_{*}^{\alpha}(B_{R}(x_{\circ}))$
$\geq\frac{1}{|B_{R}(x_{\circ})|}(1-\frac{\rho_{*}^{\alpha}(|\cdot|^{2})}{R^{2}})$
$\geq\frac{1}{|B_{R}(x_{\circ})|}(1-\frac{\rho_{0}(|\cdot|^{2})+C_{2}}{R^{2}})$
where we have used Markov’s inequality in the second inequality, and (2.7) in the last inequality. We can now take $R$ large enough such that $1-\frac{\rho_{0}(|\cdot|^{2})+C_{2}}{R^{2}}\geq\frac{1}{2}$, and obtain
$\sup\limits_{B_{R}(x_{\circ})}\rho_{*}^{\alpha}\geq\frac{1}{2|B_{R}(x_{\circ})|}\,,$
which together with the previous inequalities imply
$\rho_{*}^{\alpha}(B_{\varepsilon}(x_{\circ}))\geq\frac{|B_{\varepsilon}(x_{\circ})|}{2\,C_{B_{2R}(x_{\circ})}|B_{R}(x_{\circ})|}=:C_{\varepsilon}>0.$
3. Global convergence
We are now ready to state and prove our main results. We shall first treat the case where the function $f(\cdot)$ to be minimized has a unique minimizer. Then we expound the more general situation where $f(\cdot)$ has multiple minimizing regions.
3.1. Single–minimizer
Assumption 3.1.
$f(\cdot)$ has a unique global minimizer $x_{*}$.
The following observation is useful for the proof of the subsequent theorem.
Remark 3.2.
Thanks to the continuity and coercivity of $f(\cdot)$ in Assumption 2.1, for all arbitrary small $\varepsilon>0$ it holds that
(3.1) $\inf\{f(x):|x-x_{*}|\geq\varepsilon\}>f(x_{*})=:\underline{f}\,.$
See also Remark 3.7 for further references.
Theorem 3.3.
Assume $f(\cdot)$ satisfies Assumptions 2.1 and 3.1, and let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.4). Define the probability measure
(3.2) $\eta_{*}^{\alpha}(A):=\frac{\int_{A}e^{-\alpha f(x)}\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{*}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d}).$
Then for any $\varepsilon>0$, it holds
$\eta_{*}^{\alpha}(\{x\in{\mathbb{R}}^{d}:|x-x_{*}|\geq\varepsilon\})\longrightarrow 0\quad\mbox{ as }\quad\alpha\to\infty.$
Proof.
Let $\varepsilon>0$ be fixed, and let us define
$\overline{f}_{\varepsilon}:=\sup\{f(x):|x-x_{*}|\leq\varepsilon\}\mbox{ and }\underline{f}_{\varepsilon}:=\inf\{f(x):|x-x_{*}|\geq\varepsilon\}.$
Since $f(\cdot)$ is continuous at $x_{*}$ we know that for any $\varepsilon>0$, there exists some $\varepsilon^{\prime}>0$ such that
(3.3) $\overline{f}_{\varepsilon^{\prime}}\leq\frac{1}{2}(\underline{f}_{\varepsilon}+\underline{f}),\mbox{ that is }\quad 2(\overline{f}_{\varepsilon^{\prime}}-\underline{f})\leq(\underline{f}_{\varepsilon}-\underline{f})\,.$
Then we compute
$\eta_{*}^{\alpha}(\{x\in{\mathbb{R}}^{d}:|x-x_{*}|\geq\varepsilon\})$ $=\frac{\int_{|x-x_{*}|\geq\varepsilon}e^{-\alpha f(x)}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(x)}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}$
$\leq\frac{\int_{|x-x_{*}|\geq\varepsilon}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{|x-x_{*}|\leq\varepsilon^{\prime}}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}.$
•  In $\{x\in{\mathbb{R}}^{d}:|x-x_{*}|\geq\varepsilon\}$, we have $f(x)\geq\underline{f}_{\varepsilon}$, hence $e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\leq e^{-\alpha(\underline{f}_{\varepsilon}-\overline{f}_{\varepsilon^{\prime}})}$. Moreover, we have $\underline{f}_{\varepsilon}-\overline{f}_{\varepsilon^{\prime}}=(\underline{f}_{\varepsilon}-\underline{f})-(\overline{f}_{\varepsilon^{\prime}}-\underline{f})\geq(\overline{f}_{\varepsilon^{\prime}}-\underline{f})>0$ using (3.3). Therefore, we have
$e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\leq e^{-\alpha(\underline{f}_{\varepsilon}-\overline{f}_{\varepsilon^{\prime}})}\leq e^{-\alpha(\overline{f}_{\varepsilon^{\prime}}-\underline{f})}$
and $\rho_{*}^{\alpha}$ being a probability measure, we have
$\int_{|x-x_{*}|\geq\varepsilon}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)\leq e^{-\alpha(\overline{f}_{\varepsilon^{\prime}}-\underline{f})}\;\to 0\;\mbox{ when }\;\alpha\to\infty.$
•  In $\{x\in{\mathbb{R}}^{d}:|x-x_{*}|\leq\varepsilon^{\prime}\}$, we have $f(x)\leq\overline{f}_{\varepsilon^{\prime}}$, hence $e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\geq 1$, and
$\int_{|x-x_{*}|\leq\varepsilon^{\prime}}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)\geq\rho_{*}^{\alpha}(B_{\varepsilon^{\prime}}(x_{*}))\geq C_{\varepsilon^{\prime}}>0$
where in the last inequality $C_{\varepsilon^{\prime}}$ comes from Lemma 2.10, which is independent of $\alpha$.
Finally, one gets
$\eta_{*}^{\alpha}(\{x\in{\mathbb{R}}^{d}:|x-x_{*}|\geq\varepsilon\})$ $\leq\frac{e^{-\alpha(\overline{f}_{\varepsilon^{\prime}}-\underline{f})}}{C_{\varepsilon^{\prime}}}\,\;\to 0\;\mbox{ when }\;\alpha\to\infty.$
We can now readily state and prove the main result when the function $f(\cdot)$ to be minimized is assumed to have a unique global minimizer.
Corollary 3.4.
Let ${\overline{X}}_{.}$ be the solution to the rescaled CBO (1.4). Then in the situation of Theorem 3.3, the expectation of ${\overline{X}}_{.}$ satisfies
$\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)=\kappa\,\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})$
and
$\lim_{\alpha\to\infty}\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\kappa\,\lim_{\alpha\to\infty}\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})=\kappa\,x_{*}$
where $\rho_{*}^{\alpha}$ is the unique invariant probability measure of ${\overline{X}}_{.}$ and $x_{*}$ is the unique global minimizer of $f(\cdot)$.
Proof.
By definition, a probability measure $\rho_{*}^{\alpha}$ is invariant for ${\overline{X}}_{\cdot}$ satisfying (1.4) if and only if it is a fixed point for the adjoint of its (nonlinear) transition semigroup $\{T_{t}\}_{t\geq 0}$, that is $T^{*}_{t}\rho_{*}^{\alpha}=\rho_{*}^{\alpha}$ for all $t>0$; the adjoint operator $T^{*}_{t}$ being defined on the space of probability measures as $T^{*}_{t}\nu:=\text{Law}({\overline{X}}_{t})$ when $\nu=\text{Law}({\overline{X}}_{0})$. Recalling the dynamics (1.4), we have
(3.4) ${\overline{X}}_{t}$ $={\overline{X}}_{0}-\lambda\int_{0}^{t}({\overline{X}}_{s}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{s}))\,{\rm d}s+\sigma\int_{0}^{t}\left(\delta\,\mathds{I}_{d}+D({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho^{\alpha}_{s}))\right)\,{\rm d}B_{s}$
[…]
where $\rho^{\alpha}_{s}:=\text{Law}({\overline{X}}_{s})$. In particular, when choosing as initial distribution $\text{Law}({\overline{X}}_{0})=\rho_{*}^{\alpha}$, then it holds $\text{Law}({\overline{X}}_{t})=\rho_{*}^{\alpha}$ for all $t>0$ since the latter probability is invariant.
Let us now consider the (stationary) process (3.4) with initial distribution ${\overline{X}}_{0}\sim\rho_{*}^{\alpha}$. It is well-posed thanks to (2.9). Then taking the expectation in (3.4) and bearing in mind the invariance of $\rho_{*}^{\alpha}$, one has
$\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)=\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)-\lambda t\left(\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)-\kappa\,\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})\right)\,$
which further implies that
$\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)=\kappa\,\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha}).$
We can now consider (using again the same notation) the process (3.4) but with an arbitrary initial distribution ${\overline{X}}_{0}\sim\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$. Then, the result in (2.8) yields
$\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\lim_{t\to\infty}\;\int_{{\mathbb{R}}^{d}}x\rho_{t}^{\alpha}(\mbox{\rm d}x)=\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)$
which together with the previous equality proves the first statement of the corollary.
Next, to prove the limit when $\alpha\to\infty$, we introduce $Y_{*}^{\alpha}$ as a random variable with probability distribution $\eta_{*}^{\alpha}$ defined in (3.2). Then we have
$|\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})-x_{*}|$ $\leq{\mathbb{E}}[|Y_{*}^{\alpha}-x_{*}|]=\int_{{\mathbb{R}}^{d}}|x-x_{*}|\eta_{*}^{\alpha}(\mbox{\rm d}x)$
$=\int_{|x-x_{*}|\leq\varepsilon}|x-x_{*}|\eta_{*}^{\alpha}(\mbox{\rm d}x)+\int_{|x-x_{*}|\geq\varepsilon}|x-x_{*}|\eta_{*}^{\alpha}(\mbox{\rm d}x)$
$\leq\varepsilon+\bigg{(}\eta_{*}^{\alpha}(\{x:|x-x_{*}|\geq\varepsilon\})\bigg{)}^{\frac{1}{2}}\left(\int_{|x-x_{*}|\geq\varepsilon}|x-x_{*}|^{2}\eta_{*}^{\alpha}(\mbox{\rm d}x)\right)^{\frac{1}{2}}$
$\leq\varepsilon+\bigg{(}\eta_{*}^{\alpha}(\{x:|x-x_{*}|\geq\varepsilon\})\bigg{)}^{\frac{1}{2}}\bigg{(}2|x_{*}|^{2}+2\,\eta_{*}^{\alpha}(|\cdot|^{2})\bigg{)}^{\frac{1}{2}}$
Recalling (2.1) and (2.7), we have
(3.5) $\eta_{*}^{\alpha}(|\cdot|^{2})\leq b_{1}+b_{2}\rho_{*}^{\alpha}(|\cdot|^{2})\leq b_{1}+b_{2}(\rho_{0}(|\cdot|^{2})+C_{2})$
and then
$|\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})-x_{*}|\leq\varepsilon+\bigg{(}\eta_{*}^{\alpha}(\{x:|x-x_{*}|\geq\varepsilon\})\bigg{)}^{\frac{1}{2}}\bigg{(}2|x_{*}|^{2}+2b_{1}+2b_{2}(\rho_{0}(|\cdot|^{2})+C_{2})\bigg{)}^{\frac{1}{2}}.$
Thus, we can take $\alpha\to\infty$ while using Theorem 3.3, and obtain
$\lim_{\alpha\to\infty}\,|\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})-x_{*}|\leq\varepsilon$
which holds for any arbitrarily chosen $\varepsilon>0$. In particular, when letting $\varepsilon\to 0$, we obtain the desired result
$\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})\to x_{*}\quad\mbox{ as }\quad\alpha\to\infty.$
3.2. Multiple–minimizer
The strategy developed in the previous subsection can be further generalized to a set of minimizers that is neither a unique singleton, nor a collection of discrete points. More specifically, we shall drop the assumption on $f(\cdot)$ having a unique minimizer, and we consider the following.
Assumption 3.5.
The function $f(\cdot)$ attains its minimal value on a set $\mathcal{M}=\cup_{i=1}^{m}\,\mathcal{O}_{i}$, with $1\leq m<\infty$ and where every $\mathcal{O}_{i}$ is a compact and connected subset of ${\mathbb{R}}^{d}$.
On each $\mathcal{O}_{i}$, the function $f(\cdot)$ is constant equal to its minimal value. When $\mathcal{O}_{i}$ is not a singleton, the graph (landscape) of $f(\cdot)$ is flat, which why these are sometimes called flat minimizers. It is also worth mentioning that a particular example of Assumption 3.5 is when $f(\cdot)$ has $m$ global minimizers $\{x_{*}^{1},\dots,x_{*}^{m}\}=:\mathcal{M}$. In this case, $\mathcal{O}_{i}=\{x_{*}^{i}\}$.
Remark 3.6.
The function $f(\cdot)$ being continuous and coercive thanks to Assumption 2.1, for all arbitrary $\varepsilon>0$ it holds that
(3.6) $\inf\{f(x):x\in\mathcal{M}^{c}_{\varepsilon}\}>\underline{f}\,$
where
$\mathcal{M}_{\varepsilon}:=\cup_{i=1}^{m}\,\{x\in{\mathbb{R}}^{d}:\mathrm{dist}(x,\mathcal{O}_{i})\leq\varepsilon\}$
and $\mathcal{M}^{c}_{\varepsilon}={\mathbb{R}}^{d}\setminus\mathcal{M}_{\varepsilon}$ its complement. Here, $\mathrm{dist}(x,\mathcal{O})$ is the Euclidean distance between a point $x\in{\mathbb{R}}^{d}$ and a compact set $\mathcal{O}$, and which is equal to 00 when $x\in\mathcal{O}$.
Remark 3.7.
Property (3.6), which in the case of a single minimizer is (3.1), has already been discussed in [bardi2023eikonal, equation (H)], see also [bardi2023eikonal, Remark 2] for a counter-example.
The first result we need in the sequel is the following tightness of the weighted probability measures.
Lemma 3.8.
Let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.4), and let $\eta_{*}^{\alpha}$ be defined as in (3.2). Then $\{\eta_{*}^{\alpha}\}_{\alpha\geq 1}$ is tight.
Proof.
It follows from (2.1) and (2.7)
$\eta_{*}^{\alpha}(|\cdot|^{2})\leq b_{1}+b_{2}\,\rho_{*}^{\alpha}(|\cdot|^{2})\leq b_{1}+b_{2}\big{(}\rho_{0}(|\cdot|^{2})+C_{2}\big{)}=:C_{*}\,.$
Therefore, for any $\varepsilon$, there exists a compact set $K_{\varepsilon}:=\{x\in{\mathbb{R}}^{d}\,:\,|x|^{2}\leq\frac{C_{*}}{\varepsilon}\}$ such that, by Markov’s inequality, it holds
$\eta_{*}^{\alpha}(K_{\varepsilon}^{c})=\eta_{*}^{\alpha}\left(\left\{x\in{\mathbb{R}}^{d}\,:\,|x|^{2}>\frac{C_{*}}{\varepsilon}\right\}\right)\leq\frac{\varepsilon\,\eta_{*}^{\alpha}(|\cdot|^{2})}{C_{*}}\leq\varepsilon.$
Thanks to the latter Lemma, one applies Prokhorov’s theorem and concludes that there exist a subsequence $\{\eta_{*}^{\alpha_{k}}\}_{k\geq 1}$ and a probability measure $\eta_{*}\in\mathscr{P}({\mathbb{R}}^{d})$ such that $\eta_{*}^{\alpha_{k}}$ converges weakly to $\eta_{*}$ as $k\to\infty$. In the following we still use the notation $\eta_{*}^{\alpha}$ to denote the subsequence.
The next theorem can be viewed as a Laplace principle.
Theorem 3.9.
Let $f(\cdot)$ satisfy Assumptions 2.1 and 3.5, and let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.4). Let $\eta_{*}^{\alpha}$ be defined as in (3.2). Then, when $\alpha\to\infty$, any weak limit $\eta_{*}$ of the sequence $\{\eta_{*}^{\alpha}\}_{\alpha\geq 1}$ concentrates on $\mathcal{M}$.
Proof.
For any $\varepsilon>0$, let us define
$\overline{f}_{\varepsilon}:=\sup\{f(x):x\in\mathcal{M}_{\varepsilon}\}\mbox{ and }\underline{f}_{\varepsilon}:=\inf\{f(x):x\in\mathcal{M}_{\varepsilon}^{c}\}$
where $\mathcal{M}_{\varepsilon}$ is as defined in Remark 3.6. Since $f$ is continuous, we know that for any $\varepsilon>0$, there exists some $\varepsilon^{\prime}>0$ such that
(3.7) $\overline{f}_{\varepsilon^{\prime}}\leq\frac{1}{2}(\underline{f}_{\varepsilon}+\underline{f}),\mbox{ that is }\quad 2(\overline{f}_{\varepsilon^{\prime}}-\underline{f})\leq(\underline{f}_{\varepsilon}-\underline{f})\,.$
Then we compute
$\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})$ $=\frac{\int_{\mathcal{M}_{\varepsilon}^{c}}e^{-\alpha f(x)}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(x)}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}\;\leq\frac{\int_{\mathcal{M}_{\varepsilon}^{c}}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{\mathcal{M}_{\varepsilon^{\prime}}}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)}.$
•  In the set $\mathcal{M}_{\varepsilon}^{c}$, we have $f(x)\geq\underline{f}_{\varepsilon}$, hence $e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\leq e^{-\alpha(\underline{f}_{\varepsilon}-\overline{f}_{\varepsilon^{\prime}})}$. Moreover using (3.7), it holds
$\underline{f}_{\varepsilon}-\overline{f}_{\varepsilon^{\prime}}=(\underline{f}_{\varepsilon}-\underline{f})-(\overline{f}_{\varepsilon^{\prime}}-\underline{f})\geq(\overline{f}_{\varepsilon^{\prime}}-\underline{f})>0.$
Therefore, we have
$e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\leq e^{-\alpha(\underline{f}_{\varepsilon}-\overline{f}_{\varepsilon^{\prime}})}\leq e^{-\alpha(\overline{f}_{\varepsilon^{\prime}}-\underline{f})}$
and $\rho_{*}^{\alpha}$ being a probability measure, we have
$\int_{\mathcal{M}_{\varepsilon}^{c}}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)\leq e^{-\alpha(\overline{f}_{\varepsilon^{\prime}}-\underline{f})}\;\to 0\;\mbox{ when }\;\alpha\to\infty.$
•  In the set $\mathcal{M}_{\varepsilon^{\prime}}$, we have $f(x)\leq\overline{f}_{\varepsilon^{\prime}}$, hence $e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\geq 1$, and
$\int_{\mathcal{M}_{\varepsilon^{\prime}}}e^{-\alpha(f(x)-\overline{f}_{\varepsilon^{\prime}})}\,\rho_{*}^{\alpha}(\mbox{\rm d}x)\geq\rho_{*}^{\alpha}(\mathcal{M}_{\varepsilon^{\prime}})\geq C_{\varepsilon^{\prime}}>0$
where in the last inequality $C_{\varepsilon^{\prime}}$ comes from Lemma 2.10, which is independent of $\alpha$.
Finally, one gets
(3.8) $\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\leq\frac{e^{-\alpha(\overline{f}_{\varepsilon^{\prime}}-\underline{f})}}{C_{\varepsilon^{\prime}}}\,\;\to 0\;\mbox{ when }\;\alpha\to\infty.$
It follows from the fact that $\eta_{*}^{\alpha}$ converges weakly to $\eta_{*}$ as $\alpha\to\infty$ (Lemma 3.8), that
$\eta_{*}(A)\leq\liminf_{\alpha\to\infty}\eta_{*}^{\alpha}(A)$
for any open set $A\subset{\mathbb{R}}^{d}$. It is obvious that $\mathcal{M}_{\varepsilon}^{c}$ is an open set. Thus we have $\eta_{*}(\mathcal{M}_{\varepsilon}^{c})\leq 0\,$, which implies that $\eta_{*}(\mathcal{M}_{\varepsilon})=1\,$. Since $\varepsilon$ can be made arbitrarily small, we have indeed proved that $\eta_{*}$ concentrates on $\mathcal{M}$, that is $\eta_{*}(\mathcal{M})=1$. ∎
As a consequence of the previous theorem, we obtain a convergence in mean, hence also in probability, to the set of minimizers.
Corollary 3.10.
Let $Y_{*}^{\alpha}$ be a random variable with probability distribution $\eta_{*}^{\alpha}$ defined in (3.2). In the situation of Theorem 3.9, the following holds
$\lim\limits_{\alpha\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}=0$
where $\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})=\min\limits_{z\in\mathcal{M}}|Y_{*}^{\alpha}-z|$. In particular, it holds
$\lim\limits_{\alpha\to\infty}\,\mathbb{P}\big{(}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})=0\big{)}=1,$
Proof.
Let $\delta>0$ be fixed. Markov’s inequality yields
$\mathbb{P}\big{(}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})>\delta\big{)}\leq\frac{\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}}{\delta}.$
Then it suffices to show that $\lim\limits_{\alpha\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}=0$.
Using the definition of $Y_{*}^{\alpha}$, we have
$\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}$ $=\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We claim the following:
(3.9) $\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)$
where $L^{\infty}({\mathbb{R}}^{d},\mathcal{M})$ is the space of measurable functions from ${\mathbb{R}}^{d}$ with values in $\mathcal{M}$.
Therefore, we have
(3.10) $\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\quad\forall\,\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M}).$
We can now choose a particular function $\underline{\Delta}(\cdot)$ in the latter inequality, for example
(3.11) $\underline{\Delta}:{\mathbb{R}}^{d}\to\mathcal{M},\quad\quad\underline{\Delta}(y)=\left\{\begin{aligned} &\;\operatorname*{argmin}\limits_{z\in\mathcal{M}}|y-z|,\quad\text{if }y\in\mathcal{M}_{\varepsilon},\\ &\quad z_{*},\quad\text{if }y\in({\mathbb{R}}^{d}\setminus\mathcal{M}_{\varepsilon})=\mathcal{M}_{\varepsilon}^{c},\end{aligned}\right.$
where $z_{*}$ is some point within $\mathcal{M}$ (whose choice is without importance). Recall $\mathcal{M}_{\varepsilon}$ is as defined in Remark 3.6. Such function satisfies $|\underline{\Delta}(y)|\leq\varepsilon+|z_{*}|$. Hence we obtain
$\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\int_{{\mathbb{R}}^{d}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We split the integral into two terms accordingly
(3.12) $\int_{{\mathbb{R}}^{d}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)=\int_{\mathcal{M}_{\varepsilon}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)+\int_{\mathcal{M}_{\varepsilon}^{c}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
The integrand in the first integral is upper-bounded by $\varepsilon$ since $y\in\mathcal{M}_{\varepsilon}$, which means $|y-\underline{\Delta}(y)|\leq\varepsilon$. Thus we have
$\int_{\mathcal{M}_{\varepsilon}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\varepsilon.$
For the second integral in (3.12), we shall proceed as in the second part of the proof of Corollary 3.4. Here, we are working with $y$ in $\mathcal{M}_{\varepsilon}^{c}$, thus $\underline{\Delta}(y)=z_{*}$. Using Cauchy-Schwarz inequality, we have
(3.13) $\int_{\mathcal{M}_{\varepsilon}^{c}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)$ $\leq\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}\int_{\mathcal{M}_{\varepsilon}^{c}}|y-\underline{\Delta}(y)|^{2}\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\bigg{)}^{\frac{1}{2}}$
$\leq\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}2\,\int_{\mathcal{M}_{\varepsilon}^{c}}|y|^{2}+|\underline{\Delta}(y)|^{2}\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\bigg{)}^{\frac{1}{2}}$
$\leq\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}2\,|z_{*}|+2\,\eta_{*}^{\alpha}(|\cdot|^{2})\bigg{)}^{\frac{1}{2}}.$
Therefore, we obtain
$\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\varepsilon+\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}2\,|z_{*}|+2\,\eta_{*}^{\alpha}(|\cdot|^{2})\bigg{)}^{\frac{1}{2}}$
Recalling (2.1) and (2.7), the second moment $\eta_{*}^{\alpha}(|\cdot|^{2})$ is bounded as in (3.5).
Finally, using Theorem 3.9 (see in particular (3.8)), we know that $\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\,\to 0\,$ as $\alpha\to\infty$, hence $\lim\limits_{\alpha\to\infty}\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\varepsilon$. Here $\varepsilon$ being arbitrarily chosen, this yields $\lim\limits_{\alpha\to\infty}\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}=0$, consequently also $\lim\limits_{\alpha\to 0}\mathbb{P}\big{(}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})>\delta\big{)}=0$. The conclusion follows, recalling $\delta$ being arbitrarily chosen.
Proof of the claim (3.9):
For any $\theta>0$, there exists $\Delta_{\theta}(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})$ such that
$\int_{{\mathbb{R}}^{d}}|y-\Delta_{\theta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\theta+\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We have $\min\limits_{z\in\mathcal{M}}|y-z|\leq|y-\Delta_{\theta}(y)|$. Integrating with respect to $\eta_{*}^{\alpha}$ yields
$\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\theta+\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
This holds for any arbitrarily chosen $\theta$, in particular when $\theta\to 0$, which then proves the claim and concludes the proof. ∎
Remark 3.11.
It can be shown in fact that the claim (3.9) holds with an equality instead of an inequality. The proof of the other inequality relies on a selection principle; see [kouhkouh2024viscous, Proposition 3.1].
Corollary 3.12.
Let $Y_{t}^{\alpha}$ be a random variable whose probability distribution is $\eta_{t}^{\alpha}$ defined by
(3.14) $\eta_{t}^{\alpha}(A):=\frac{\int_{A}e^{-\alpha f(x)}\rho_{t}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{t}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d}),\quad\rho_{t}=\mathrm{Law}({\overline{X}}_{t}).$
Then in the situation of Theorem 3.9, we have
$\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}=0.$
In particular, we have
$\lim\limits_{\alpha\to\infty}\lim\limits_{t\to\infty}\,\mathbb{P}\big{(}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})=0\big{)}=1.$
Note that the latter convergence holds independently of the initial distribution of $Y_{0}^{\alpha}\sim\eta_{0}^{\alpha}$ defined with $\rho_{0}=\mathrm{Law}({\overline{X}}_{0})\in\mathscr{P}_{4}({\mathbb{R}}^{d})$.
Proof.
Let $\delta>0$ be fixed. Markov’s inequality yields
$\mathbb{P}\big{(}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})>\delta\big{)}\leq\frac{\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}}{\delta}.$
Then it suffices to show that $\lim\limits_{\alpha\to\infty}\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}=0$.
Using the definition of $Y_{t}^{\alpha}$, we have
$\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}$ $=\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{t}^{\alpha}(\mbox{\rm d}y)$
$\leq\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{t}^{\alpha}(\mbox{\rm d}y)$
where the inequality follows from (3.9).
Letting $\alpha$ be fixed, we introduce the notation:
$F_{t}:L^{\infty}({\mathbb{R}}^{d},\mathcal{M})\to\mathbb{R},\quad\quad F_{t}(\Delta):=\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{t}^{\alpha}(\mbox{\rm d}y)$
and, analogously, we define
$F_{*}:L^{\infty}({\mathbb{R}}^{d},\mathcal{M})\to\mathbb{R},\quad\quad F_{*}(\Delta):=\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We shall first prove that the sequence $\{F_{t}\}_{t>0}$ converges uniformly to $F_{*}$ when $t\to\infty$. Consequently we can exchange the minimization and the limit $t\to\infty$, and conclude as in the proof of Corollary 3.10.
Step 1. (Uniform convergence)
Recalling the definition of $\eta_{t}^{\alpha}$ in (3.14), we can express $F_{t}$ such that
$F_{t}(\Delta)=\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho^{\alpha}_{t}(\mbox{\rm d}y)$
where $h_{\Delta}(y)=|y-\Delta(y)|e^{-\alpha f(y)}\quad\text{ and }\quad g(\rho^{\alpha}_{t})=\int_{{\mathbb{R}}^{d}}e^{-\alpha f(x)}\,\rho_{t}^{\alpha}(\mbox{\rm d}x).$
Analogously, we can also write
$F_{*}(\Delta)=\frac{1}{g(\rho_{*}^{\alpha})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y).$
Thus we have
$|F_{t}(\Delta)-F_{*}(\Delta)|$ $\leq\left|\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho^{\alpha}_{t}(\mbox{\rm d}y)-\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)\right|$
$\quad\quad\quad+\left|\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)-\frac{1}{g(\rho_{*}^{\alpha})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)\right|$
$\leq\frac{1}{g(\rho^{\alpha}_{t})}\left|\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho^{\alpha}_{t}(\mbox{\rm d}y)-\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)\right|$
$\quad\quad\quad+\left|\frac{1}{g(\rho^{\alpha}_{t})}-\frac{1}{g(\rho_{*}^{\alpha})}\right|\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y).$
The set $\mathcal{M}$ being compact, one can find a constant $C>0$ such that for all $\Delta\in L^{\infty}(\mathbb{R}^{d},\mathcal{M})$ we have $|h_{\Delta}(y)|\leq C(1+|y|)$. Hence, using (2.8) (see also [zhang2023existence, Lemma 2.7]), we can send $t\to\infty$ in each term of the latter inequality and obtain the desired uniform convergence.
Step 2. (Exchange of minimization and limit)
Consider a quasi-minimizer of $F_{*}$, that is, for any $\varepsilon>0$, there exists $\Delta_{\varepsilon}$ such that
(3.15) $F_{*}(\Delta_{\varepsilon})\leq\varepsilon+\min\limits_{\Delta}F_{*}(\Delta).$
The uniform convergence guarantees that for any $\varepsilon>0$, there exists $T_{\varepsilon}>0$ such that for all $t\geq T_{\varepsilon}$ it holds
(3.16) $|F_{t}(\Delta)-F_{*}(\Delta)|\leq\varepsilon\quad\forall\,\Delta.$
In particular, when the latter is evaluated in $\Delta_{\varepsilon}$, one has
$F_{t}(\Delta_{\varepsilon})\leq F_{*}(\Delta_{\varepsilon})+\varepsilon$
which together with (3.15) yields
$F_{t}(\Delta_{\varepsilon})\leq 2\,\varepsilon+\min\limits_{\Delta}F_{*}(\Delta).$
But $\min\limits_{\Delta}F_{t}(\Delta)\leq F_{t}(\Delta_{\varepsilon})$. Therefore we have
$\min\limits_{\Delta}F_{t}(\Delta)\leq 2\,\varepsilon+\min\limits_{\Delta}F_{*}(\Delta).$
Taking first the limit $t\to\infty$, then letting $\varepsilon\to 0$, yield $\lim\limits_{t\to\infty}\min\limits_{\Delta}F_{t}(\Delta)\leq\min\limits_{\Delta}F_{*}(\Delta)$.
We now prove the converse inequality. Given $t>0$, let us consider $\{\Delta_{n,t}\}_{n\in\mathbb{N}}$ a minimizing sequence for $F_{t}(\cdot)$, that is
(3.17) $\lim\limits_{n\to\infty}F_{t}(\Delta_{n,t})=\min\limits_{\Delta}F_{t}(\Delta).$
The inequality (3.16) evaluated in $\Delta_{n,t}$ yields (when choosing $t\geq T_{\varepsilon}$)
$F_{*}(\Delta_{n,t})\leq\varepsilon+F_{t}(\Delta_{n,t}).$
But $\min\limits_{\Delta}F_{*}(\Delta)\leq F_{*}(\Delta_{n,t})$. Therefore we have
$\min\limits_{\Delta}F_{*}(\Delta)\leq\varepsilon+F_{t}(\Delta_{n,t}).$
We can now first take the limit $n\to\infty$ using (3.17), then we let $t\to\infty$, which yield $\min\limits_{\Delta}F_{*}(\Delta)\leq\varepsilon+\lim\limits_{t\to\infty}\min\limits_{\Delta}F_{t}(\Delta)$. The conclusion follows by sending $\varepsilon\to 0$.
Step 3. (Conclusion)
We have shown that
$\lim\limits_{t\to\infty}\,\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}F_{t}(\Delta)=\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}F_{*}(\Delta)$
which then yields
$\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}$ $\leq\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)$
$\leq\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\quad\quad\forall\,\Delta\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M}).$
In particular we can evaluate the latter inequality in $\underline{\Delta}$ defined in (3.11), and conclude as in Corollary 3.10 that
$\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}=0.$
4. Conclusion
We have studied asymptotic properties of the rescaled CBO dynamics
${\rm d}{\overline{X}}_{t}=-\lambda\left({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}^{\alpha})\right)\,{\rm d}t+\sigma\left(\delta\,\mathds{I}_{d}+D({\overline{X}}_{t}-\kappa\,\mathfrak{m}_{\alpha}(\rho_{t}^{\alpha}))\right){\rm d}B_{t},\quad\rho_{t}^{\alpha}=\text{Law}({\overline{X}}_{t})$
under suitable conditions on the parameters, for example
$\delta=1,\quad\kappa<\frac{1}{2}(1+b_{2})^{-1},\quad\text{ and }\quad\lambda>16\,\sigma^{2}$
where $b_{2}$ is a constant depending on the function $f(\cdot)$ to be minimized (see Assumption 2.1 and Lemma 2.2). We do not make any assumptions on the initial distribution ${\overline{X}}_{0}\sim\rho_{0}$ besides satisfying $\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$.
Recalling (2.8) obtained from [zhang2023existence, Lemma 2.7], for any $|\varphi(x)|\leq C(1+|x|)$ it holds
$\lim_{t\to\infty}\int_{{\mathbb{R}}^{d}}\varphi(x)\,\mbox{\rm d}\rho_{t}^{\alpha}(x)=\int_{{\mathbb{R}}^{d}}\varphi(x)\,\mbox{\rm d}\rho_{*}^{\alpha}(x)\,$
where $\rho_{*}^{\alpha}$ is the unique invariant probability measure of the rescaled CBO. When combined with (3.4), we obtain
(4.1) $\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\kappa^{-1}{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\kappa^{-1}\int_{{\mathbb{R}}^{d}}x\,\mbox{\rm d}\rho_{t}^{\alpha}(x)\,=\,x_{*}$
where $x_{*}$ is the unique global minimizer of $f(\cdot)$.
When $f(\cdot)$ has multiple minimizers, we obtain the concentration of the weighted probability measures
$\eta_{*}^{\alpha}(A)=\frac{\int_{A}e^{-\alpha f(x)}\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{*}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d}).$
on the set of global minimizers, in the limit $\alpha\to\infty$. Moreover, if we let $Y_{*}^{\alpha}$ be a random variable with probability distribution $\eta_{*}^{\alpha}$, and $Y_{t}^{\alpha}$ be a random variable whose probability distribution is
$\eta_{t}^{\alpha}(A)=\frac{\int_{A}e^{-\alpha f(x)}\rho_{t}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{t}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d}),\quad\rho_{t}=\mathrm{Law}\left({\overline{X}}_{t}\right),$
then we have shown that it converges in mean, hence also in probability, to the set of global minimizers $\mathcal{M}$. More precisely we have
(4.2) $\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}=\lim\limits_{\alpha\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}=0,$
for all initial distributions $Y_{0}^{\alpha}\sim\eta_{0}^{\alpha}$ given with $\rho_{0}=\mathrm{Law}\left({\overline{X}}_{0}\right)\in\mathscr{P}_{4}({\mathbb{R}}^{d})$.